Sam has to eat one and one half pies in two minutes at the end of one minute he ate seven-eighths of a pie. How much pie does Sam have left to eat?

Answer by mananth(16946)

1 1/2 pie = 3/2 pie
he ate 7/8
so balance to be eaten
= (3/2)-(7/8)
=(12-7)/8
= 5/8 ia balance
...
m.ananth@hotmail.ca